# Easy attention: A simple attention mechanism for temporal predictions with transformers

### v3

@@ Introduction
Dynamical systems are mathematical models used to describe the evolution of complex phenomena over time. These systems are prevalent in various fields, including physics and engineering. A dynamical system consists of variables that change over time, and their interactions are governed by differential equations. The behavior of such systems can range from simple and predictable to highly complex and chaotic. Furthermore, chaotic systems are characterized by their sensitivity to perturbations of the initial conditions, leading to unpredictability and complex dynamics. The prediction of chaotic dynamical systems stands as a relevant and challenging area of study. Accurate forecasting of such systems holds significant importance in diverse scientific and engineering disciplines, including weather forecasting, ecological modeling, financial analysis, and control of complex physical processes.
@@
Dynamical-system models leveraging the concept of time delays [1, 2, 3, 4] have been applied for temporal-dynamics prediction. These models also leverage decomposition to obtain a reduced-order representation of the system. For instance, proper-orthogonal decomposition [5] provides a set of deterministic spatial functions with corresponding time coefficients. The decomposition can also be performed in the spectral domain [6]. Another widely used method in different scientific fields is the dynamic-mode decomposition (DMD) [7, 8] which captures intricate temporal patterns and can reveal underlying modes of motion. DMD represents a numerical approach to the approximation of the Koopman operator [2, 9, 10, 11]. Koopman-based frameworks have achieved promising results in the prediction of temporal dynamics of high-dimensional chaotic systems [12, 13, 14, 15, 16]. In the context of time-series analysis, the well-known family of auto-regressive and moving-average (ARMA) models is used for forecasting dynamical systems [17]. These methods focus on stochastic problems and they yield interpretable models [4]. Deep neural networks (DNNs), as a branch of machine-learning (ML) approaches, have demonstrated remarkable success in temporal-dynamics prediction, offering advantages of flexibility and adaptability. The family of recurrent neural networks (RNNs) [18] has been the dominant method for temporal-dynamics prediction for many years, enabling the model to capture temporal dependencies in the time series. Among RNNs, the long short-term memory (LSTM) networks [19] overcome the training problems related to vanishing gradients and this enables modelling long-term dependencies. Herefore, LSTM networks are effective ML models for temporal predictions of complex dynamical systems [20, 21, 16]. Besides RNNs, the temporal convolutional networks (TCNs) [22], which employ one-dimensional convolutions for multi-variable time-series prediction [23, 24], have shown promising performance on temporal-dynamics prediction [25].
@@
In recent years, transformer neural networks [26] have revolutionized the field of ML, demonstrating remarkable capabilities in natural-language processing (NLP) [27] and computer vision [28], but also fluid mechanics [29]. Thanks to their matrix-multiplication-based attention mechanism [26], transformers have shown the capability to identify and predict the temporal dynamics of chaotic systems by capturing long-term dependencies in the data [25, 30, 31]. This promising performance has motivated several studies focused on improving the attention mechanism of transformer models used for physical modeling [32, 33]. Additionally, in the context of reducing complexity in NLP, more studies on improving attention mechanism, in particular self-attention, have emerged [34, 35, 36], and they have reported promising results while reducing the computational cost. Note that all the mentioned improvements on self-attention follow the idea of the retrieval process which comprises queries, keys and values. However, in the present study, inspired by work done on Koopman operators [13, 14, 15, 16], we propose a new attention mechanism denoted by easy attention, which use neither the $\rm softmax$ [32, 35] nor the linear projections of queries and keys; this has the goal of reducing the complexity of the model and improving its interpretability. Our objective is to predict the temporal evolution of dynamical systems. Typically, self-attention is computed for every input; however, this choice can be less effective in chaotic systems as the input can vary arbitrarily, while keeping the same underlying system dynamics. Conservative attractors remain invariant under small perturbations as proposed by Kolmogorov–Arnold–Moser (KAM) Theory [37]. Based on this intuition, we propose an easy-attention formulation that is optimized during training and it is not input-dependent at inference time. Doing so, we learn an importance score for the input variables that encodes the system dynamics rather than the instantaneous attention of the input values. This approach provides significant performance improvements in the prediction of the trajectories of chaotic systems, as it fully exploits the spatial periodicity that is naturally present in most chaotic systems, assuming an underlying constant pattern over time, the invariant set.
Interpretation of attention mechanism
@@
In this section, we provide a motivating example for simplifying the self-attention mechanism and propose the easy-attention technique described in the Methods section. We start with a reconstruction task of a simple wave function, using the self-attention module. Based on the data, we investigate the eigenspace given by the self-attention mechanism, and analyze the main limitation of this mechanism. Subsequently, we discuss how to address such limitation using the easy attention method and we compare the reconstruction results.
@@
where $t$ is time spanning the range $[0,3N]$ and each $y_{i}$ a trajectory. A satisfactory performance of the different attention mechanisms for this case is a prerequisite to attempt the prediction of more challenging problems. Here our focus is on identification and reconstruction of the wave motion and we assess the results obtained just with the attention module, without any extra hidden layers and activation functions. For this case, each sample consists of the three sine waves with different phase shifts at 3 consecutive time instants, $\mathbf{y_{i}}=\{y_{i}(t_{3p}-k\Delta t)\}_{k=0}^{2}$, where $\Delta t=1$. Interpreting $\mathbf{y}_{i}$ as a column vector, we can write the input matrix for the attention mechanism as $\mathbf{Y}=\text{concat}\{{\bf y}_{i}\}_{i=1}^{3}$, where concat refers to concatenation, as represented in Eq. (2). Here we refer to $p$ as the current sample in the batch and $n$ as the time delay.
@@
Note that the query and key matrices are used to compute the dot product $\langle\mathbf{Q}(\mathbf{y}_{i}),\mathbf{K}(\mathbf{y}_{j})\rangle$ and the attention scores $\alpha_{ij}={\rm softmax}_{j}(\langle\mathbf{Q}(\mathbf{y}_{i}),\mathbf{K}(\mathbf{y}_{j})\rangle/\sqrt{k})$, where $k$ is the well known dmodel size for transformers or latent dimension. Since the trace of the operator is invariant to cyclic permutations of its argument, we can write:
@@
This enables identifying two matrices: $\mathbf{W}_{Q}\mathbf{W}_{K}^{T}$ and $\mathbf{y_{j}}\mathbf{y_{i}^{T}}$. We will refer to $\alpha_{ij}$ as the transformed covariance factor given that $\bm{\alpha}=\mathbf{C}\iff\mathbf{W_{Q}}\mathbf{W_{K}}^{T}=\mathbb{I}$, where $\mathbb{I}$ is the identity matrix, $\mathbf{C}=\mathbf{Y}^{T}\mathbf{Y}$ is the covariance matrix of the input matrix $\mathbf{Y}$ and $C_{i,j}=\rm{Tr}(\mathbf{y_{j}}\mathbf{y_{i}}^{T})$ the covariance factor; note that $\mathbf{y}_{j}\mathbf{y}_{i}^{T}$ is the covariance factor matrix. Due to the rank-1 nature of the covariance factor matrix, performing the eigen-decomposition of the latter would not allow us to extract any relevant information. In order to address this issue, we apply singular-value decomposition (SVD) [38] to the matrix product $\mathbf{W_{Q}}\mathbf{W_{K}}^{T}\mathbf{y}_{j}\mathbf{y}_{i}^{T}$ in order to gain insight into its structure.
@@
The SVD of the transformed covariance factor matrix given by $\mathbf{y}_{j}\mathbf{y}_{i}^{T}$ as well as the product of the query and key matrices used to later compute the attention scores, together with the corresponding scores in $\bm{\alpha}$ are shown in Fig. 1. Note that we only show the main diagonal elements of the $\bm{\alpha}$ matrix as these are the main contributors when computing the attention score. From the transformed covariance factor matrix, one can observe that the spanned space by the left singular vectors, $\mathbf{u}$, exhibits an approximately cylindrical shape. However, the right singular vectors ${\bm{\phi}}$ exhibit a shape closer to that of a sphere. These shapes are connected within the unit circle, which is used to study unique sinusoidal functions. When analyzing multiple waves with different phase shifts, given orthogonal phase shifts, the unit circle is expected to grow in dimensions as the transformed covariance factor matrix between input vectors increases. It can be claimed that as $r\rightarrow\infty$ for $\mathbf{Y}\in\mathbb{R}^{3\times r}$, the sphere will become denser, creating a fully compact surface if and only if the new waves in Eq. (2) are linearly independent with the rest of input columns. The eigenspaces for queries and keys, shown in Fig. 1(b), are observed to be fixed by rotations and translations of $\mathbf{W}_{Q}\mathbf{W}_{K}^{T}$. These orbits constitute the subspace of the real domain, where the inner product exists. However, after the space compression forced by the trace and $\rm softmax$, this subdomain is reduced to a scalar, which is shown in Fig. 1(c) as a non-trivial space compression. One can interpret $\bm{\alpha}$ as the transformed covariance matrix between inputs of the transformer, such that $\alpha_{i,j}\in[0,1],\forall i,j$. In fact, the $\rm{softmax}$ is a probability distribution over three different possible trajectories, each trajectory being an input vector. For instance, the second element in the first row of $\bm{\alpha}$ is expressed as:
@@
In Eq. (6), it can be observed that the numerator fulfills $\exp\big{(}{\rm Tr}(\mathbf{W}_{Q}\mathbf{W}_{K}^{T}\mathbf{y}_{2}\mathbf{y}_{1}^{T})/\sqrt{k}\big{)}=\exp\big{(}\sum_{i=1}^{3}\lambda_{i}/\sqrt{k}\big{)}$, where $\lambda_{i}$ represent the eigenvalues of the matrix, and this is, after normalization, the attention factor between $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$: $\alpha_{1,2}$. The latter behaves like a probability distribution: when we combine all possible trajectories spanned by all possible $x_{j}=\mathbf{V(y_{j})}$ with their respective attention factors, we obtain the expected trajectory as a reconstruction for the next time instances of the time series $\mathbf{y}_{1}$, expressed as $\mathbf{\hat{y}}_{1}$. Therefore, for a fixed $i$ and $X_{i}=\{x_{1},x_{2},x_{3}\}$ with probabilities $p_{j}=\alpha_{i,j}$:
@@
It is important to recognize the difference between language (for which the self-attention framework was mainly developed) and physical problems. When analyzing text, the attention matrix is mostly diagonal, nearly a bijective mapping, as for each word that we input there exists a (near) one-to-one relationship with another word in the ground truth. However, for physical problems this behavior cannot be assumed and it this becomes obvious during the training phase.
Based on the discussion above, the nature of the trace eliminates all backwards traceability of the information, decreasing the relevance of the exact value of each matrix element. Therefore, it is reasonable to directly learn the exact $\bm{\alpha}$ value instead of learning the query and key matrices. Here, we propose removing the $\mathbf{W}_{K}$, $\mathbf{W}_{Q}$ and $\rm softmax$ from the self attention, which (assuming square matrices $\mathbf{W}_{K}$ and $\mathbf{W}_{Q}$ after embedding into the latent dimension $k$) reduces the computational complexity of the problem from $4n_{i}^{2}$ to $2n_{i}^{2}$, where $n_{i}$ is the size of the input vectors, $\mathbf{y}_{i}(t)$.
@@
At this stage, we employ the easy-attention module using an input of the same size as that of the self attention, and the same training setup for sinusoidal-waves reconstruction. We evaluate the reconstruction accuracy by the relative $l_{2}$-norm error defined as:
@@
where $S$ is the ground-truth sequence and $\tilde{S}$ is the prediction obtained by the attention module. In the Supplementary Material we show details regarding $l_{2}$-norm errors, computation time for training ($t_{c}$) and number of floating-point operations ($N_{f}$) of the employed easy-attention and self-attention modules. The results indicate that easy attention reduces the complexity of self attention by 45%, saves computation time by 15% and outperforms the self-attention, with an error of $0.0018\%$ compared with the $10\%$ given by the latter.
Temporal-dynamics predictions of chaotic systems using a transformer network with easy attention
In the following section we will use a modified transformer encoder with easy attention in order to conduct a systematic study regarding the performance of the latter against other state-of-the-art deep-learning models. In particular, we first study the Lorenz system (a three-dimensional canonical chaotic system), secondly the Moehlis model [40] for wall-bounded turbulence (which is a system comprising nine ordinary differential equations) and finally a model [41] of a Fluoride-salt-cooled high-temperature (FHR) reactor.

### v4

@@ I Introduction
Dynamical systems are mathematical models used to describe the evolution of complex phenomena over time. These systems are prevalent in various fields, including physics and engineering. A dynamical system consists of variables that change over time, and their interactions are governed by differential equations. The behavior of such systems can range from simple and predictable to highly complex and chaotic. Furthermore, chaotic systems are characterized by their sensitivity to perturbations of the initial conditions, leading to unpredictability and complex dynamics [1]. The prediction of chaotic dynamical systems stands as a relevant and challenging area of study. Accurate forecasting of such systems holds significant importance in diverse scientific and engineering disciplines, including weather forecasting, ecological modeling, financial analysis, and control of complex physical processes.
@@
Dynamical-system models leveraging the concept of time delays [2, 3, 4, 5] have been applied for temporal-dynamics prediction. These models also leverage decomposition to obtain a reduced-order representation of the system. For instance, proper-orthogonal decomposition [6] provides a set of deterministic spatial functions with corresponding time coefficients. The decomposition can also be performed in the spectral domain [7]. Another widely used method in different scientific fields is the dynamic-mode decomposition (DMD) [8, 9] which captures intricate temporal patterns and can reveal underlying modes of motion. DMD represents a numerical approach to the approximation of the Koopman operator [3, 10, 11, 12]. Koopman-based frameworks have achieved promising results in the prediction of temporal dynamics of high-dimensional chaotic systems [13, 14, 15, 16, 17]. In the context of time-series analysis, the well-known family of auto-regressive and moving-average (ARMA) models is used for forecasting dynamical systems [18]. These methods focus on stochastic problems and they yield interpretable models [5]. Deep neural networks (DNNs), as a branch of machine-learning (ML) approaches, have demonstrated remarkable success in temporal-dynamics prediction, offering advantages of flexibility and adaptability [19]. The family of recurrent neural networks (RNNs) [20] has been the dominant method for temporal-dynamics prediction for many years, enabling the model to capture temporal dependencies in the time series. Among RNNs, the long short-term memory (LSTM) networks [21] overcome the training problems related to vanishing gradients and this enables modelling long-term dependencies. Herefore, LSTM networks are effective ML models for temporal predictions of complex dynamical systems [22, 23, 17]. Besides RNNs, the temporal convolutional networks (TCNs) [24], which employ one-dimensional convolutions for multi-variable time-series prediction [25, 26], have shown promising performance on temporal-dynamics prediction [27].
@@
In recent years, transformer neural networks [28] have revolutionized the field of ML, demonstrating remarkable capabilities in natural-language-processing (NLP) [29] and computer vision [30], but also fluid mechanics [31]. Thanks to their matrix-multiplication-based attention mechanism [28], transformers have shown the capability to identify and predict the temporal dynamics of chaotic systems by capturing long-term dependencies in the data [27, 32, 33]. This promising performance has motivated several studies focused on improving the attention mechanism of transformer models used for physical modeling [34, 35]. Furthermore, in the context of reducing computational complexity in natural language processing tasks such as machine translation and language modeling, more studies have emerged on improving the attention mechanism, in particular self-attention, have emerged [36, 37, 38], and they have reported promising results while reducing computational cost. Note that all the mentioned improvements on self-attention follow the idea of the retrieval process which comprises queries, keys and values. However, in the present study, the approach is inspired by recent work on Koopman operators, which suggests treating attention as a finite-rank operator [14, 15, 16, 17], we propose a new attention mechanism denoted by easy attention, which uses neither the $\rm softmax$ [34, 37] nor the linear projections of queries and keys; this has the goal of reducing the complexity of the model and improving its interpretability. Our objective is to predict the temporal evolution of dynamical systems. Typically, self-attention is computed for every input; however, this choice can be less effective in chaotic systems as the input can vary arbitrarily, while keeping the same underlying system dynamics. The Kolmogorov–Arnold–Moser (KAM) Theory shows that conservative attractors remain invariant under small perturbations [39]. Based on this intuition, we propose an easy-attention formulation that is optimized during training and it is not input-dependent at inference time. Thus, we learn an importance score for the input variables that encodes the system dynamics rather than the instantaneous attention of the input values. This approach provides significant performance improvements in the prediction of the trajectories of chaotic systems, as it fully exploits the spatial periodicity that is naturally present in most chaotic systems, assuming an underlying constant pattern over time, the invariant set.
This paper is structured as follows: Section II exploits the self-attention mechanism performance eigendecomposing the $\alpha$ to elaborate the rationality of easy-attention. Section III discusses the application of easy-attention in a transformer architecture for temporal-dynamic prediction, and shows the results. The conclusions are presented in Section IV and finally, Section V offers an overview of self-attention mechanisms and introduces the easy-attention mechanism in detail.
II Interpretation of attention mechanism
@@
In this section, we provide a motivating example for simplifying the self-attention mechanism to the proposed easy-attention technique described in the Methods section. We start with a reconstruction task of a simple wave function, using the self-attention module. Based on the data, we investigate the eigenspace given by the self-attention mechanism, and analyze the main limitation of this mechanism. Subsequently, we discuss how to address such limitation using the easy attention method and we compare the reconstruction results.
@@
where $t$ is time spanning the range $[0,3N]$ and each $y_{i}$ a trajectory. A satisfactory performance of the different attention mechanisms for this case is a prerequisite to attempt the prediction of more challenging problems. Here our focus is on identification and reconstruction of the wave motion and we assess the results obtained just with the attention module, without any extra hidden layers and activation functions. For this case, each sample consists of the three sine waves with different phase shifts at 3 consecutive time instants, $\mathbf{y_{i}}=\{y_{i}(t_{3p}-k\Delta t)\}_{k=0}^{2}$, where $\Delta t=1$. Interpreting $\mathbf{y}_{i}$ as a column vector, we can write the input matrix for the attention mechanism as $\mathbf{Y}=\text{concat}\{{\bf y}_{i}\}_{i=1}^{3}$, where concat refers to concatenation, as represented in Eq. (2). Here we refer to $p$ as the current sample in the batch and $k\Delta t$ as the time delay.
@@
Note that the query and key matrices are used to compute the dot product $\langle\mathbf{Q}(\mathbf{y}_{i}),\mathbf{K}(\mathbf{y}_{j})\rangle$ and the attention scores $\alpha_{ij}={\rm softmax}_{j}(\langle\mathbf{Q}(\mathbf{y}_{i}),\mathbf{K}(\mathbf{y}_{j})\rangle/\sqrt{k})$, where $k$ is the well known $\rm{d_{model}}$ size for transformers or latent dimension. Since the trace of the operator is invariant to cyclic permutations of its argument, we can write:
@@
where $\operatorname{Tr}$ denotes the sum of diagonal entries of its matrix argument (i.e. the usual matrix trace). Intuitively, $\operatorname{Tr}(\mathbf{M})$ measures the total covariance captured along each coordinate direction of the vectors feeding into $\mathbf{M}$, where $\mathbf{M}=\mathbf{W}_{Q}\mathbf{W}_{K}^{T}\mathbf{y}_{j}\mathbf{y}_{i}^{T})$.This enables identifying two matrices: $\mathbf{W}_{Q}\mathbf{W}_{K}^{T}$ and $\mathbf{y_{j}}\mathbf{y_{i}^{T}}$. We will refer to $\alpha_{ij}$ as the transformed covariance factor given that $\bm{\alpha}=\mathbf{C}\iff\mathbf{W_{Q}}\mathbf{W_{K}}^{T}=\mathbb{I}$, where $\mathbb{I}$ is the identity matrix, $\mathbf{C}=\mathbf{Y}^{T}\mathbf{Y}$ is the covariance matrix of the input matrix $\mathbf{Y}$ and $C_{i,j}=\rm{Tr}(\mathbf{y_{j}}\mathbf{y_{i}}^{T})$ the covariance factor; note that $\mathbf{y}_{j}\mathbf{y}_{i}^{T}$ is the covariance factor matrix. In order to gain insight into its structure, we apply singular-value decomposition (SVD) [40] to the matrix product $\mathbf{W_{Q}}\mathbf{W_{K}}^{T}\mathbf{y}_{j}\mathbf{y}_{i}^{T}$.
@@
The SVD of the transformed covariance factor matrix given by $\mathbf{y}_{j}\mathbf{y}_{i}^{T}$ as well as the product of the query and key matrices used to later compute the attention scores, together with the corresponding scores in $\bm{\alpha}$ are shown in Fig. 1. Note that we only show the main diagonal elements of the $\bm{\alpha}$ matrix as these are the main contributors when computing the attention score. From the transformed covariance factor matrix, one can observe that the spanned space by the left singular vectors, $\mathbf{u}$, exhibits an approximately cylindrical shape. However, the right singular vectors ${\bm{\phi}}$ exhibit a shape closer to that of a sphere. These shapes are connected within the unit circle, which is used to study unique sinusoidal functions. When analyzing multiple waves with different phase shifts, the unit circle is expected to grow in dimensions as the transformed covariance factor matrix between input vectors increases. It can be claimed that as $r\rightarrow\infty$ for $\mathbf{Y}\in\mathbb{R}^{3\times r}$; $r$ being the size of the time-delays, the sphere will become denser, creating a fully compact surface if and only if the new waves in Eq. (2) are linearly independent with the rest of input columns. The eigenspaces for queries and keys, shown in Fig. 1(b), are observed to be fixed by rotations and translations of $\mathbf{W}_{Q}\mathbf{W}_{K}^{T}$. These orbits constitute the subspace of the real domain, where the inner product exists. However, after the space compression forced by the trace and $\rm softmax$, this subdomain is reduced to a scalar, which is shown in Fig. 1(c) as a non-trivial space compression. One can interpret $\bm{\alpha}$ as the transformed covariance matrix between inputs of the transformer, such that $\alpha_{i,j}\in[0,1],\forall i,j$, where $\alpha_{i,j}$ is defined as the scalar entry in row i, column j of the matrix $\mathbf{\alpha}$. In fact, the $\rm{softmax}$ is a probability distribution over three different possible trajectories, each trajectory being an input vector. For instance, the second element in the first row of $\bm{\alpha}$ is expressed as:
@@
In Eq. (6), it can be observed that the numerator fulfills $\exp\big{(}{\rm Tr}(\mathbf{W}_{Q}\mathbf{W}_{K}^{T}\mathbf{y}_{2}\mathbf{y}_{1}^{T})/\sqrt{k}\big{)}=\exp\big{(}\sum_{i=1}^{3}\lambda_{i}/\sqrt{k}\big{)}$, where $\lambda_{i}$ represent the eigenvalues of the matrix product, and this is, after normalization, the attention factor between $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$: $\alpha_{1,2}$. The latter is a probability distribution: when we combine all possible trajectories spanned by all possible $x_{j}=\mathbf{V(y_{j})}$ with their respective attention factors, we obtain the expected trajectory as a reconstruction for the next time instances of the time series $\mathbf{y}_{1}$, expressed as $\mathbf{\hat{y}}_{1}$. Therefore, for a fixed $i$ and $X_{i}=\{x_{1},x_{2},x_{3}\}$ with probabilities $p_{j}=\alpha_{i,j}$:
@@
It is important to recognize the difference between language (for which the self-attention framework was mainly developed) and physical problems. When analyzing text, the attention matrix is mostly diagonal, nearly a bijective mapping, as for each word that we input there exists a (near) one-to-one relationship with another word in the ground truth. However, for physical problems this behavior cannot be assumed and this becomes obvious during the training phase.
Based on the discussion above, it is reasonable to directly learn the exact $\bm{\alpha}$ value instead of learning the query and key matrices. Here, we propose removing the $\mathbf{W}_{K}$, $\mathbf{W}_{Q}$ and $\rm softmax$ from the self attention, which (assuming square matrices $\mathbf{W}_{K}$ and $\mathbf{W}_{Q}$ after embedding into the latent dimension $k$) reduces the computational complexity of the problem from $4n_{i}^{2}$ to $2n_{i}^{2}$, where $n_{i}$ is the size of the input vectors, $\mathbf{y}_{i}(t)$.
@@
At this stage, we employ the easy-attention approach using an input of the same size as that of the self attention, and the same training setup for sinusoidal-waves reconstruction. We evaluate the reconstruction accuracy by the relative $l_{2}$-norm error defined as:
@@
where $S$ is the ground-truth sequence and $\tilde{S}$ is the prediction obtained by the attention module. In the Supplementary Material we show details regarding $l_{2}$-norm errors, computation time for training ($t_{c}$) and number of floating-point operations ($N_{f}$) of the employed easy-attention and self-attention modules. The results indicate that easy attention reduces complexity by 50%, as compared to self-attention, under identical embedding dimensions, saves computation time by 26% and outperforms the self-attention, with an error of $0.0018\%$ compared with the $10\%$ given by the latter. To conclude the intepretation of the attention mechanism,
III Temporal-dynamics predictions of chaotic systems using a transformer
In the following section we will use a modified transformer encoder with easy attention in order to conduct a systematic study regarding the performance of the latter against other state-of-the-art deep-learning models, LSTM and self-attention transformers. In particular, we first study the Lorenz system (a three-dimensional canonical chaotic system), secondly the Moehlis model [42] for wall-bounded turbulence (which is a system comprising nine ordinary differential equations) and finally a model [43] of a Fluoride-salt-cooled high-temperature (FHR) reactor.